# How to pick the best anomaly detector?

### v1

@@ I Introduction
In recent years there has been an explosion of new methods for model-agnostic new physics searches at the LHC, leveraging powerful modern machine-learning (ML) anomaly detection methods such as autoencoders and weak supervision [Kasieczka:2021xcg, Aarrestad:2021oeb, Belis:2023mqs, hepmllivingreview]. These methods are able to find anomalies in the data with varying degrees of model independence: either signal-model-independence, in that they make no use of any specific new physics models in their training, and/or background-model-independence in that they make no use of Standard Model (SM) simulations in their training. As such, these new model-agnostic search strategies offer significant new opportunities for discovery in the vast, unexplored phase space of the LHC data, beyond what model-specific search strategies have been able to achieve to date.
@@
So far, when it comes to different approaches (e.g. autoencoders vs. weak supervision), the idea has been to just try them all, as seen in Ref. [CMS:2024nsz]. Within a given approach, there are myriad smaller but still vital choices to make: which model epoch to pick during training, which model architecture and hyperparameters to use, or even which features to use. Metrics traditionally used to guide these choices, such as the SIC and area under curve (AUC), are based on truth labels and require the use of benchmark signals. This is clearly not satisfactory for anomaly detection as the added signal dependence may, in the worst case, work against finding a signal actually present in the data.
Consequently, none of the existing experimental anomaly detection analyses [ATLAS:2020iwa, ATLAS:2023azi, ATLAS:2023ixc, CMS:2024nsz, ATLAS:2025obc, ATLAS:2025kuz, CMS:2025dbd, CMS:2025eou] report systematic model optimization. Model architectures and training parameters are generally derived from the original publication introducing the corresponding method, with adjustments to network capacity to deal with the larger experimental datasets. In some cases [ATLAS:2023ixc, CMS:2024nsz, CMS:2025dbd], parameters have been retuned using a small set of benchmark signals. In Refs. [CMS:2024nsz, CMS:2025dbd], anomaly scores have been selected from different classifier runs or epochs in certain approaches – CWoLa Hunting and TNT used the number of events selected in the signal region at a fixed background efficiency in the side bands as a metric; and CATHODE used the validation loss, following Ref. [Hallin:2021wme].
@@
where $\epsilon_{\text{SR/BT}}$ is the efficiency to select SR/BT events for a given anomaly threshold. In contrast to SIC, which cannot be calculated for real (unlabeled) data, ARGOS is accessible for any anomaly detector if there is a corresponding background template in addition to the (unlabeled) data set. In Section II, we show that ARGOS has a sound theoretical foundation, being monotonic with the actual signal-to-background SIC for any unknown signal at a fixed background efficiency when the background template is perfect. More generally, ARGOS is proportional to SIC to a very good approximation in any meaningful setting.
For concreteness, we demonstrate the utility of ARGOS within the context of weak supervision [Metodiev:2017vrx], where a classifier is trained to distinguish two mixed-label data sets, for resonant anomaly detection. However, nothing in principle limits ARGOS to this context. Using the LHCO dataset [Kasieczka:2021xcg], introduced in Section III, we demonstrate how ARGOS can consistently select the best hyperparameters, architecture (e.g. neural networks vs. different boosted decision trees) and model epoch in Section IV. As an outlook, we also briefly discuss the potential for ARGOS to select the best features in a fully data-driven way before we conclude in Section V. Technical aspects of our experiments are discussed in several Appendices.
@@ II The ARGOS Metric
  • At fixed $\epsilon_{B}$, ARGOS is monotonic with SIC, since the signal fraction $f_{S}$ is a constant. Then the best anomaly detector with respect to the ARGOS metric is also the best anomaly detector with respect to SIC. Additionally, ARGOS and the number of selected signal region events are also monotonic in this case and therefore also result in the same optimal decisions. The latter was used in Ref. [CMS:2024nsz, CMS:2025dbd] to select epochs from multiple trainings of the CWoLa Hunting and TNT approaches.
  • For ARGOS, however, there is no need to fix the background efficiency a priori. For a good anomaly detector, $\sqrt{\epsilon_{B}}\ll{\rm SIC}$ also holds, such that the maximum value of ARGOS as a function of score threshold is also close to the maximum value of SIC. Then an added benefit of using the ARGOS metric in this way is that one can also select the optimal working point of the anomaly detector in a data-driven way.
@@ III.1 Data set
In Section IV, we perform experiments in a number of different scenarios, which test different metrics for optimizing the setup for anomaly detection. These experiments are performed using the LHC Olympics 2020 (LHCO) [Kasieczka:2021xcg] R&D data set [LHCOdataset], which is a common benchmark data set for LHC anomaly detection. It contains $10^{6}$ QCD dijet background events and $10^{5}$ signal events. The signal is a $W^{\prime}$ resonance at $m_{W^{\prime}}=3.5\,\text{TeV}$ with the decay $W^{\prime}\rightarrow X(\rightarrow qq)Y(\rightarrow qq)$, where $X$ and $Y$ have masses $m_{X}=100\,\text{GeV}$ and $m_{Y}=500\,\text{GeV}$, respectively. Additionally, we use 612858 SR background events from Ref. [extraLHCOdataset], which were generated for Ref. [Hallin:2021wme].
The simulation pipeline consists of Pythia 8 [Sjostrand:2006za, Sjostrand:2007gs] and Delphes 3.4.1 [deFavereau:2013fsa]. Using Fastjet [Cacciari:2011ma], reconstructed particles are clustered with the anti-$k_{T}$ algorithm [Cacciari:2005hq] with a distance parameter of $R=1$. All events must pass a leading-jet trigger using $p_{\mathrm{T}}>$1.2\text{\,}\mathrm{TeV}$\$. For the classification, we use the features also used in Ref. [Hallin:2021wme]. We select the two highest $p_{T}$ jets and use the mass of the lighter jet $m_{J_{1}}$, the mass differences of the two jets $\Delta m_{J}=m_{J_{2}}-m_{J_{1}}$ as well as both jets’ 21-subjettiness ratios [Thaler:2010tr, Thaler:2011gf]. In Sec. IV.5, additional feature sets based on Ref. [Finke:2023ltw] are used. The dijet mass $m_{JJ}$ is used as the resonant feature.
@@ III.2 Weakly Supervised Methods for Resonant Anomaly Detection
We focus on classifier-based resonant anomaly detection in this work, since these methods are predicated on the existence of an accurate BT that describes the smooth background in the SR. For resonant anomalies, the data-driven construction of the BT is usually based on adjacent sideband (SB) regions that facilitate smooth interpolation into the SR. There are a variety of different methods of constructing the background template [Metodiev:2017vrx, Collins:2018epr, Collins:2019jip, Nachman:2020lpy, Andreassen:2020nkr, 1815227, Hallin:2021wme, Raine:2022hht, Hallin:2022eoq, Golling:2022nkl, Golling:2023yjq, Das:2024fwo, Leigh:2024chm, Oleksiyuk:2025pmu], and we use the following three to illustrate the utility of the ARGOS metric:
@@ III.2.1 Idealized Anomaly Detector
The idealized anomaly detector (IAD) [Hallin:2021wme] is used as a benchmark which is available for simulated data only. Instead of constructing the BT in a data-driven way, the IAD simply uses background events simulated with the same pipeline that generated the background data in the SR. Hence, by definition, $p^{\text{IAD}}_{\text{BT}}(x)=p^{\text{ideal}}_{\text{BT}}(x)=p_{B}(x)$. We use approximately 272 000272\,000 such events.
@@ III.2.2 CWoLa Hunting
CWoLa Hunting [Collins:2018epr, Collins:2019jip] directly uses data from short sidebands of $0.2\,\text{TeV}$ on either side of the signal region as the background template, assuming there is little correlation of the features with the dijet invariant mass. Using short sidebands reduces correlations with the dijet mass. As these short sidebands are in total as wide as the SR, the resulting background template has a similar size to the SR.
@@ III.2.3 CATHODE
CATHODE [Hallin:2021wme] uses conditional density estimation to interpolate the SB distribution into the SR: as the SB contains only very low amounts of signal events, learning $p(x|m\in SB)$ is largely equivalent to learning the background distribution $p_{B}(x|m)$, which can then be sampled in the SR to obtain background events in the SR. The density estimator can be used to generate more BT events than there are SR events in order to obtain better background statistics. We use a fixed oversampling factor of four as in Ref. [CMS:2024nsz]. The implementation of the density estimation is discussed in App. A.1.
@@ III.3 Classifiers
    A standard multi-layer perceptron (MLP) neural network (NN) with three hidden layers, which is trained using the binary cross-entropy (BCE) loss with the Adam optimizer for 100 epochs. We ensemble the best 10 epochs according to our different metrics, following the ensembling strategy used in Ref. [Hallin:2021wme], where the best 10 epochs based on the validation loss are used.
@@
    The HistGradientBoostingClassifier, which is a gradient boosted decision tree classifier implementation in scikit-learn [Pedregosa:2011sk]. It histograms the classification features before training. We ensemble 50 independent trainings with different 50-50 training-validation splits as was done in Ref. [Finke:2023ltw].
@@
    The AdaBoostingClassifier, which is an adaptive boosted decision tree classifier implementation in scikit-learn [Pedregosa:2011sk]. We build our implementation in analogy to Ref. [Finke:2023ltw] and perform an ensembling of 10 independently trained classifiers using random 50-50 splits of the training dataset.
@@ III.4 Performance Metrics
    The BCE loss calculated on the validation set, which is the standard metric used to perform model optimization or select the best epochs during training (see e.g. Refs. [Hallin:2021wme, CMS:2024nsz]).
@@ IV.2 NN epoch selection
An interesting feature can be seen in the performance of the max SIC based epoch selection at $N_{sig}=0$, where the performance is slightly better than random. The fluctuations of the classifier weights in the absence of signal still lead to some epochs with non-zero anomaly-detection performance, which the max SIC identifies. Selecting the epochs using signal information can therefore enhance the false positive rate. Using ARGOS, which does not know anything about a potential signal, this issue is not present.
@@ IV.5 Feature selection
For anomaly detection to be truly signal model agnostic, feature sets should not be tuned on specific signals. To test the concept of choosing the best feature set based on data-driven metrics, we apply the procedure used for the architecture selection (see Sec. IV.4) to the feature sets used in Ref. [Finke:2023ltw] for the IAD in Fig. 6. The largest feature set, extended set 3, is in this case most sensitive and at high signal injections always chosen by the ARGOS metric. At lower signal injections ARGOS sometimes also selects less performing feature sets as can be seen from the increased error band and the slightly earlier performance drop-off.
Figure 6: Anomaly detection performance max SIC for different subjettiness-based feature sets and for the feature set selection using max ARGOS in black. We use the IAD setup with the HGB classifier. The baseline set is defined in Section III. The extended sets are introduced in Ref. [Finke:2023ltw].
@@ V Conclusion and Outlook
We stress that this is not the only context where the ARGOS metric can be used. For example, it is directly applicable to ANODE [Nachman:2020lpy] and R-ANODE [Das:2023bcj], which are density-estimator-based resonant anomaly detection methods. Beyond resonant anomaly detection, different outlier detectors (e.g. autoencoders) could also be compared to one another if a background template is available. In contrast to the binary cross-entropy loss for classifier-based anomaly detection, ARGOS does not assume predictions to represent a class probability. ARGOS is a cut-based metric, which is only affected by the relative ordering of different events and invariant to rescaling or shifting of the entire anomaly score. Therefore, ARGOS can be applied toany anomaly score.
@@ A.1 Implementation of the CATHODE Density Estimation
In order to obtain state of the art results, we use a conditional flow matching (CFM) density estimator based on Ref. [Das:2024fwo]. In this implementation [density_estimation], the conditional vector field needed for CFM is learned using a ResNet-style [resnet] architecture from nflows [nflows]. The hyperparameters from Ref. [Das:2024fwo] are used for both the model and the training. For each signal injection, ten density estimators are trained and ensembled by combining their samples in order to limit fluctuations of the density estimation quality between the different signal injections.
@@ A.2 Classifier implementations and hyperparameter settings
For the NN, we use a standard MLP based on the architecture used in Ref. [Hallin:2021wme] with an implementation in PyTorch [Ansel_PyTorch_2_Faster_2024] based on Ref. [sk-cathode]. We choose a fixed number of three hidden layers with 64 nodes each and ReLU activation. We train on the binary cross-entropy loss using the ADAM optimizer [Kingma:2014ad] for 100 epochs. Other hyperparameters such as batch size, dropout, learning rate and other ADAM settings are optimized later on. Their default values can be found in Tab. 1. For the NN, we draw hyperparameters from a set of choices, which are also included in the table.
We choose a 50-50 training and validation split, calculating BCE and max ARGOS on the validation set and max SIC on the test set after every epoch. After training, we ensemble the best 10 epochs as decided by the different metrics. This follows the ensembling strategy used in Ref. [Hallin:2021wme], where the best 10 epochs based on the validation loss are used. The impact of choosing different metrics can be seen in Sec. IV.2. As a default, we pick the epochs on max ARGOS.

### v2

@@ I Introduction
In recent years there has been an explosion of new methods for model-agnostic new physics searches at the LHC, leveraging powerful modern machine-learning (ML) anomaly detection methods such as autoencoders and weak supervision [22, 2, 6, 21]. These methods are able to find anomalies in the data with varying degrees of model independence: either signal-model-independence, in that they make no use of any specific new physics models in their training, and/or background-model-independence in that they make no use of Standard Model (SM) simulations in their training. As such, these new model-agnostic search strategies offer significant new opportunities for discovery in the vast, unexplored phase space of the LHC data, beyond what model-specific search strategies have been able to achieve to date.
@@
So far, when it comes to different approaches (e.g. autoencoders vs. weak supervision), the idea has been to just try them all, as seen in Ref. [8]. Within a given approach, there are myriad smaller but still vital choices to make: which model epoch to pick during training, which model architecture and hyperparameters to use, or even which features to use. Metrics traditionally used to guide these c qhoices, such as the SIC and area under curve (AUC), are based on truth labels and require the use of benchmark signals. This is clearly not satisfactory for anomaly detection as the added signal dependence may, in the worst case, work against finding a signal actually present in the data.
Consequently, none of the existing experimental anomaly detection analyses [ATLAS:2020iwa, 4, ATLAS:2023ixc, 8, 1, 5, CMS:2025dbd, CMS:2025eou] report systematic model optimization. Model architectures and training parameters are generally derived from the original publication introducing the corresponding method, with adjustments to network capacity to deal with the larger experimental datasets. In some cases [ATLAS:2023ixc, 8, CMS:2025dbd], parameters have been retuned using a small set of benchmark signals. In Refs. [8, CMS:2025dbd], anomaly scores have been selected from different classifier runs or epochs in certain approaches – CWoLa Hunting and TNT used the number of events selected in the signal region at a fixed background efficiency in the side bands as a metric; and CATHODE used the validation loss, following Ref. [19]. In Ref. [17], combining different anomaly detection models has been studied in the context of the NLPM methodology [11].
@@
where $\epsilon_{\text{SR/BT}}$ is the efficiency to select SR/BT events for a given anomaly threshold. By definition, ARGOS is equivalent to the standard test statistic for an excess in a cut-and-count experiment with $N_{\text{SR}}$ signal region events, where the observed number of events $\epsilon_{\text{SR}}N_{\text{SR}}$ after the anomaly cut is compared to the expected event number $\epsilon_{\text{BT}}N_{\text{SR}}$ for the null hypothesis. This test statistic has also been used for specific anomaly detection methods [Das:2024eie, 18]. Here, however, we advocate ARGOS as a metric for optimizing anomaly detection performance. In contrast to SIC, which cannot be calculated for real (unlabeled) data, ARGOS is accessible for any anomaly detector if there is a corresponding background template in addition to the (unlabeled) data set. In Section II, we show that ARGOS has a sound theoretical foundation, being monotonic with the actual signal-to-background SIC for any unknown signal at a fixed background efficiency when the background template is perfect. More generally, ARGOS is proportional to SIC to a very good approximation in any meaningful setting.
For concreteness, we demonstrate the utility of ARGOS within the context of weak supervision [24], where a classifier is trained to distinguish two mixed-label data sets, for resonant anomaly detection. However, nothing in principle limits ARGOS to this context. Using the LHCO dataset [22], introduced in Section III, we demonstrate how ARGOS can consistently select the best hyperparameters, architecture (e.g. neural networks vs. different boosted decision trees) and model epoch in Section IV. As an outlook, we also briefly discuss the potential for ARGOS to select the best features in a fully data-driven way before we conclude in Section V. Technical aspects of our experiments are discussed in several Appendices.
@@ II The ARGOS Metric
  • At fixed $\epsilon_{B}$, ARGOS is monotonic with SIC, since the signal fraction $f_{S}$ is a constant. Then the best anomaly detector with respect to the ARGOS metric is also the best anomaly detector with respect to SIC. Additionally, ARGOS and the number of selected signal region events are also monotonic in this case and therefore also result in the same optimal decisions. The latter was used in Ref. [8, CMS:2025dbd] to select epochs from multiple trainings of the CWoLa Hunting and TNT approaches.
  • For ARGOS, however, there is no need to fix the background efficiency a priori. For a good anomaly detector, $\sqrt{\epsilon_{B}}\ll{\rm SIC}$ also holds, such that the maximum value of ARGOS as a function of score threshold is also close to the maximum value of SIC. Then an added benefit of using the ARGOS metric in this way is that one can also select the optimal working point of the anomaly detector in a data-driven way. In Ref. [18], instead anomaly detectors with different background efficiencies have been combined using the corresponding test statistic.
@@ III.1 Data set
In Section IV, we perform experiments in a number of different scenarios, which test different metrics for optimizing the setup for anomaly detection. These experiments are performed using the LHC Olympics 2020 (LHCO) [22] R&D data set [LHCOdataset], which is a common benchmark data set for LHC anomaly detection. It contains $10^{6}$ QCD dijet background events and $10^{5}$ signal events. The signal is a $W^{\prime}$ resonance at $m_{W^{\prime}}=3.5\,\text{TeV}$ with the decay $W^{\prime}\rightarrow X(\rightarrow qq)Y(\rightarrow qq)$, where $X$ and $Y$ have masses $m_{X}=100\,\text{GeV}$ and $m_{Y}=500\,\text{GeV}$, respectively. Additionally, we use 612858 SR background events from Ref. [extraLHCOdataset], which were generated for Ref. [19].
The simulation pipeline consists of Pythia 8 [Sjostrand:2006za, Sjostrand:2007gs] and Delphes 3.4.1 [deFavereau:2013fsa]. Using Fastjet [Cacciari:2011ma], reconstructed particles are clustered with the anti-$k_{T}$ algorithm [Cacciari:2005hq] with a distance parameter of $R=1$. All events must pass a leading-jet trigger using $p_{\mathrm{T}}>$1.2\text{\,}\mathrm{TeV}$\$. For the classification, we use the features also used in Ref. [19]. We select the two highest $p_{T}$ jets and use the mass of the lighter jet $m_{J_{1}}$, the mass differences of the two jets $\Delta m_{J}=m_{J_{2}}-m_{J_{1}}$ as well as both jets’ 21-subjettiness ratios [Thaler:2010tr, Thaler:2011gf]. In Sec. IV.5, additional feature sets based on Ref. [13] are used. The dijet mass $m_{JJ}$ is used as the resonant feature.
@@ III.2 Weakly Supervised Methods for Resonant Anomaly Detection
We focus on classifier-based resonant anomaly detection in this work, since these methods are predicated on the existence of an accurate BT that describes the smooth background in the SR. For resonant anomalies, the data-driven construction of the BT is usually based on adjacent sideband (SB) regions that facilitate smooth interpolation into the SR. There are a variety of different methods of constructing the background template [24, 9, 10, 25, 3, 7, 19, 27, 20, Golling:2022nkl, 16, 12, 23, 26], and we use the following three to illustrate the utility of the ARGOS metric:
@@ III.2.1 Idealized Anomaly Detector
The idealized anomaly detector (IAD) [19] is used as a benchmark which is available for simulated data only. Instead of constructing the BT in a data-driven way, the IAD simply uses background events simulated with the same pipeline that generated the background data in the SR. Hence, by definition, $p^{\text{IAD}}_{\text{BT}}(x)=p^{\text{ideal}}_{\text{BT}}(x)=p_{B}(x)$. We use approximately 272 000272\,000 such events.
@@ III.2.2 CWoLa Hunting
CWoLa Hunting [9, 10] directly uses data from short sidebands of $0.2\,\text{TeV}$ on either side of the signal region as the background template, assuming there is little correlation of the features with the dijet invariant mass. Using short sidebands reduces correlations with the dijet mass. As these short sidebands are in total as wide as the SR, the resulting background template has a similar size to the SR.
@@ III.2.3 CATHODE
CATHODE [19] uses conditional density estimation to interpolate the SB distribution into the SR: as the SB contains only very low amounts of signal events, learning $p(x|m\in SB)$ is largely equivalent to learning the background distribution $p_{B}(x|m)$, which can then be sampled in the SR to obtain background events in the SR. The density estimator can be used to generate more BT events than there are SR events in order to obtain better background statistics. We use a fixed oversampling factor of four as in Ref. [8]. The implementation of the density estimation is discussed in App. A.1.
@@ III.3 Classifiers
    A standard multi-layer perceptron (MLP) neural network (NN) with three hidden layers, which is trained using the binary cross-entropy (BCE) loss with the Adam optimizer for 100 epochs. We ensemble the best 10 epochs according to our different metrics, following the ensembling strategy used in Ref. [19], where the best 10 epochs based on the validation loss are used.
@@
    The HistGradientBoostingClassifier, which is a gradient boosted decision tree classifier implementation in scikit-learn [Pedregosa:2011sk]. It histograms the classification features before training. We ensemble 50 independent trainings with different 50-50 training-validation splits as was done in Ref. [13].
@@
    The AdaBoostingClassifier, which is an adaptive boosted decision tree classifier implementation in scikit-learn [Pedregosa:2011sk]. We build our implementation in analogy to Ref. [13] and perform an ensembling of 10 independently trained classifiers using random 50-50 splits of the training dataset.
@@ III.4 Performance Metrics
    The BCE loss calculated on the validation set, which is the standard metric used to perform model optimization or select the best epochs during training (see e.g. Refs. [19, 8]).
@@ IV.2 NN epoch selection
An interesting feature can be seen in the performance of the max SIC based epoch selection at $N_{sig}=0$, where the performance is slightly better than random. The fluctuations of the classifier weights in the absence of signal still lead to some epochs with non-zero anomaly-detection performance, which the max SIC identifies. Selecting the epochs using signal information can therefore enhance the false positive rate. Using ARGOS, which does not know anything about a potential signal, this issue is not present. We discuss why we do not generally expect an enhanced rate of false discoveries or enhanced background sculpting when selecting models on data in App. D.
@@ IV.5 Feature selection
For anomaly detection to be truly signal model agnostic, feature sets should not be tuned on specific signals. To test the concept of choosing the best feature set based on data-driven metrics, we apply the procedure used for the architecture selection (see Sec. IV.4) to the feature sets used in Ref. [13] for the IAD in Fig. 6. The largest feature set, extended set 3, is in this case most sensitive and at high signal injections always chosen by the ARGOS metric. At lower signal injections ARGOS sometimes also selects less performing feature sets as can be seen from the increased error band and the slightly earlier performance drop-off.
Figure 6: Anomaly detection performance max SIC for different subjettiness-based feature sets and for the feature set selection using max ARGOS in black. We use the IAD setup with the HGB classifier. The baseline set is defined in Section III. The extended sets are introduced in Ref. [13].
@@ V Conclusion and Outlook
We stress that this is not the only context where the ARGOS metric can be used. For example, it is directly applicable to ANODE [25] and R-ANODE [Das:2023bcj], which are density-estimator-based resonant anomaly detection methods. Beyond resonant anomaly detection, different outlier detectors (e.g. autoencoders) could also be compared to one another if a background template is available. In contrast to the binary cross-entropy loss for classifier-based anomaly detection, ARGOS does not assume predictions to represent a class probability. ARGOS is a cut-based metric, which is only affected by the relative ordering of different events and invariant to rescaling or shifting of the entire anomaly score. Therefore, ARGOS can be applied to any anomaly score.
@@ A.1 Implementation of the CATHODE Density Estimation
In order to obtain state of the art results, we use a conditional flow matching (CFM) density estimator based on Ref. [12]. In this implementation [density_estimation], the conditional vector field needed for CFM is learned using a ResNet-style [resnet] architecture from nflows [nflows]. The hyperparameters from Ref. [12] are used for both the model and the training. For each signal injection, ten density estimators are trained and ensembled by combining their samples in order to limit fluctuations of the density estimation quality between the different signal injections.
@@ A.2 Classifier implementations and hyperparameter settings
For the NN, we use a standard MLP based on the architecture used in Ref. [19] with an implementation in PyTorch [Ansel_PyTorch_2_Faster_2024] based on Ref. [sk-cathode]. We choose a fixed number of three hidden layers with 64 nodes each and ReLU activation. We train on the binary cross-entropy loss using the ADAM optimizer [Kingma:2014ad] for 100 epochs. Other hyperparameters such as batch size, dropout, learning rate and other ADAM settings are optimized later on. Their default values can be found in Tab. 1. For the NN, we draw hyperparameters from a set of choices, which are also included in the table.
We choose a 50-50 training and validation split, calculating BCE and max ARGOS on the validation set and max SIC on the test set after every epoch. After training, we ensemble the best 10 epochs as decided by the different metrics. This follows the ensembling strategy used in Ref. [19], where the best 10 epochs based on the validation loss are used. The impact of choosing different metrics can be seen in Sec. IV.2. As a default, we pick the epochs on max ARGOS.